a train moves from rest to a speed of 25 m/a in 30.0 seconds. what is the magnitude of its acceleration?
1
Expert's answer
2020-12-04T07:32:58-0500

We can find the magnitude of the acceleration of the train from the kinematic equation:


"v=v_0+at,""a=\dfrac{v-v_0}{t}=\dfrac{25\ \dfrac{m}{s}-0\ \dfrac{m}{s}}{30\ s}=0.83\ \dfrac{m}{s^2}."

Answer:

"a=0.83\ \dfrac{m}{s^2}."